OpenStudy (anonymous):

The grades on the last math exam had a mean of 88%. Assume the population of grades on math exams is known to be normally distributed with a standard deviation of 6. What percent of students earn a score between 72% and 90%? 0.6255 0.5843 0.6754 0.5346

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@mtbender74

OpenStudy (kirbykirby):

Let \(X\) be the grade of math exams. Since \(X\) is normally distributed, you can standardize it to find the probability asked. The question is asking \(P(72<X<90)\). Let's standardize: \[ P(72<X<90)=P\left( \frac{72-88}{6}<\frac{X-88}{6}<\frac{90-88}{6}\right)=P(-2.67<Z<0.33)\] You will then need to find the area under the standard normal distribution between -2.67 and 0.33